A Complete Bibliography of Scientific American: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

01 March 2018 Version 4.05

Title word cross-reference

1/f [?].

10th [?]. 16th [?]. 16th-Century [?].

9 [?].

Abortion [?]. Abrasive [?]. Absolute [?]. Absorption [?, ?, ?]. Accelerator [?]. Accelerators [?, ?, ?]. Accuracy [?]. Acetylcholine [?]. Acid [?]. Acoustic [?, ?]. Acoustics [?]. Acquires [?]. Acquisition [?]. Act [?]. Actinomycin [?]. Action [?, ?]. Activated [?, ?]. Active [?]. Activity [?]. Adaptation [?]. Additives [?]. Adjacency [?]. Advanced [?]. Advances [?]. Adversary [?]. Advice [?, ?]. Advice-Taking [?, ?]. Aerodynamic [?]. aether [?]. Afar [?]. Africa [?]. African [?]. Aftereffects [?]. Against [?, ?]. Age [?, ?, ?, ?, ?]. Ago [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Ago [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Agreement [?]. Agricultural [?, ?, ?, ?, ?]. Agriculture [?, ?, ?, ?, ?, ?, ?, ?]. Air [?, ?, ?]. Airborne [?, ?]. Aircraft [?, ?]. Airflow [?]. Airglow [?]. Airplane [?]. Akhenaten [?]. Albatrosses [?]. Albinos [?]. Alcohol [?]. Alfred [?]. Algae [?]. Algorithms [?, ?]. Alienation [?]. Alkali [?]. Alloying [?]. Alloys [?, ?]. Alternative [?]. Altitude [?]. Amateur [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Ambiguity [?]. America [?]. American [?]. Amorphous [?, ?]. Amorphous-Semiconductor [?, ?]. AMP [?]. Amplification [?]. Analysis [?, ?, ?, ?, ?, ?, ?]. Anatomy [?]. Anchovy [?]. Ancient [?, ?]. Andes [?, ?]. Andreas [?]. Andromeda [?]. Animal [?, ?, ?, ?, ?]. Animals [?, ?, ?, ?, ?, ?]. Anions [?]. Annihilation [?, ?]. Annual [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Antibodies [?]. Antibody [?]. Antigens [?]. Antisubmarine [?]. Ants [?, ?, ?, ?]. Ape [?]. Apollo [?]. Appendicularians [?]. Approach [?, ?]. Archaeology [?, ?]. Archaic [?, ?]. Architecture [?, ?]. Arithmetic [?]. Arms [?, ?]. Army [?]. Aromatic [?, ?]. Arrow [?]. Artificial [?]. Assembly [?, ?, ?, ?]. Assessment [?, ?]. Assistance [?]. Asteroids [?]. Astounding [?]. Astrolabe [?]. Astronomy [?]. Astrophysics [?, ?]. Asymmetry [?]. Atherosclerosis [?]. Atlantic [?, ?, ?]. Atmosphere [?, ?]. Atmospheric [?, ?]. Atomic [?]. Atoms [?, ?]. ATP [?]. Attack [?]. Attitudes [?, ?]. Attractant [?]. Auditory [?, ?]. Australia [?]. Authors [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Authors [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Automata [?]. Automatic [?, ?]. Automobile [?, ?]. Automobiles [?, ?]. Autonomic [?]. Available [?].

Background [?, ?]. Bacteria [?, ?, ?, ?, ?]. Bacterial [?]. Bag [?]. Ballistic [?]. Ban [?, ?]. Bantu [?]. Barr [?]. Based [?]. Basis [?, ?]. Batavia [?]. Beam [?, ?]. Beams [?, ?]. Bear [?]. Bearing [?]. Bearings [?]. Beats [?]. Beehive [?]. Bees [?]. Beetle [?]. Beetles [?]. before [?]. Began [?]. Behavior [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. between [?, ?, ?, ?, ?, ?, ?, ?, ?]. Bibliography [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?].

Bibliography [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Bicycle [?]. Binary [?]. binding [?]. Binds [?]. Binocular [?, ?]. Biocrystals [?]. Biological [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Biology [?]. Biosphere [?, ?, ?, ?, ?, ?]. Bird [?, ?]. Birds [?, ?, ?, ?, ?]. Birth [?, ?]. BL [?]. Black [?, ?, ?, ?]. Blacks [?]. Blindness [?]. Blood [?, ?]. Blue [?]. Blue-Green [?]. Bodies [?, ?]. Body [?, ?]. Bok [?]. Bomb [?]. Bomber [?]. Bones [?]. Book [?]. Books [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Books [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Bottom [?]. Bound [?]. Bowed [?]. Bracing [?]. Brain [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Brain-Cooling [?, ?]. Brains [?]. Brasses [?]. Braun [?]. break [?]. Breakers [?]. Breakup [?]. Breathe [?, ?]. Breeder [?, ?, ?]. Brideprice [?]. Brightest [?]. Brines [?]. Britain [?, ?, ?]. Brother [?]. brown [?]. Bruegel [?]. Bruno [?]. Bubbles [?, ?]. Building [?, ?]. Buildings [?]. Bumblebee [?]. Burgess [?]. Burial [?]. Bursts [?, ?]. Burying [?]. Business [?]. Byzantine [?].

C [?]. Cable [?]. Calcitonin [?]. Calculator [?]. Calefaction [?]. Can [?]. Canals [?]. Cancer [?, ?, ?, ?, ?]. Candidates [?]. Capture [?]. Car [?]. Carbenes [?]. Carbon [?, ?, ?, ?, ?, ?]. Carcinogens [?]. Care [?, ?, ?, ?, ?]. Carnivorous [?]. Carroll [?]. Carrying [?]. Carthage [?]. Carthaginian [?]. Case [?]. Castings [?]. Casts [?]. Catalysis [?]. Catapults [?]. Cataract [?]. Catastrophe [?]. Cathedrals [?]. Cathode [?]. Cats [?]. Cause [?]. Causes [?, ?]. Cave [?]. Cell [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Cell-Surface [?]. Cells [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Cellular [?, ?]. Celtic [?]. Cement [?]. Cemetery [?]. Census [?]. Center [?, ?, ?, ?]. Center-Pivot [?]. central [?]. Centrifuge [?]. Century [?]. Cerebellum [?]. Ceremonial [?]. Chain [?]. Chains [?]. Change [?]. Changes [?]. Changing [?]. Channels [?, ?]. Characteristics [?]. Charge [?, ?]. Charge-Coupled [?, ?]. Charm [?]. Checks [?]. Chemical [?, ?, ?, ?, ?, ?, ?]. Chemicals [?]. Chemistry [?, ?, ?, ?, ?]. Chess [?, ?]. Child [?]. Children [?]. Chimeras [?]. Chimpanzee [?]. China [?, ?, ?]. Chinese [?, ?]. Chloroplasts [?]. Choice [?]. Cholera [?]. Christ [?]. Chromosomal [?]. Chromosomes [?]. Cilia [?]. cipher [?]. Circuit [?, ?]. Circuits [?, ?, ?]. Circulation [?, ?]. Citizen [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Citizen [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. City [?]. Clams [?]. Class [?, ?, ?]. Clay [?]. Clean [?, ?, ?]. Climate [?, ?]. Clock [?, ?]. Clocks [?, ?]. Cluster [?, ?]. clustering [?]. Clusters [?, ?, ?]. Coal [?, ?, ?, ?, ?]. Coatings [?]. Code [?]. Coil [?]. Colicins [?]. Collective [?, ?]. Collective-Effect [?, ?]. Colliculus [?]. collision [?]. Collisions [?, ?]. Color [?, ?, ?, ?, ?]. Columbian [?]. Combinations [?]. Combinatorial [?]. Combining [?]. Comet [?]. Comets [?]. Commerce [?]. Communication [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Communications [?, ?, ?]. Community [?, ?]. Companions [?]. Comparisons [?]. Compartments [?]. Complement [?]. Complexity [?, ?]. Composite [?]. Compound [?]. Compounds [?]. Computer [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Computer-controlled [?, ?]. Computer-Managed [?, ?]. Computers [?]. Computing [?]. Concepts [?]. Conduction [?]. Conductors [?]. Confederacy [?]. Confinement [?, ?]. Confirmation [?]. Confusions [?]. Consequences [?]. Constants [?]. Constructive [?]. Consumer [?]. Consumption [?]. Continent [?]. Continent-Building [?]. Continental [?, ?]. Continents [?, ?]. Contour [?]. Contours [?]. Contraction [?, ?]. Contrast [?, ?]. Control [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. controlled [?, ?]. Controversies [?]. Convection [?]. Conversion [?, ?]. Conway [?]. Cooling [?, ?, ?, ?]. Cooperative [?]. Coordination [?]. Copernican [?]. Copernicus [?]. Coprolites [?]. Coral [?]. Corals [?]. Corn [?]. Corona [?]. Cortex [?]. Cosmic [?, ?, ?, ?, ?, ?, ?, ?]. Counterfeiting [?]. Countries [?, ?, ?, ?, ?, ?, ?]. Coupled [?, ?, ?]. Covenanters [?]. Crashworthiness [?]. Cratering [?, ?]. Cricket [?]. Crisis [?, ?]. Criteria [?]. Critters [?]. Crocodile [?]. Crocodilians [?]. Crops [?]. Cruise [?]. Crust [?]. Cry [?]. Cryptobiosis [?]. Cryptography [?, ?]. Crystal [?, ?, ?]. Crystals [?, ?]. Culture [?]. Currents [?, ?, ?]. Curvature [?]. curves [?]. Cutting [?]. Cyanate [?]. Cycle [?, ?, ?, ?, ?, ?, ?, ?, ?]. Cycles [?, ?, ?, ?]. Cyclic [?].

Data [?]. Dating [?, ?]. Day [?]. Daylight [?]. Daylight-Saving [?]. Death [?]. Debate [?, ?, ?]. Decay [?]. Decision [?]. Decision-making [?]. Deep [?, ?, ?]. Deep-Sea [?]. Defense [?]. Deformation [?, ?]. Deimos [?]. Deinstitutionalization [?]. Delayed [?]. Delivery [?, ?, ?]. Deposits [?, ?]. Deprivation [?]. Depth [?]. Desert [?]. Deserted [?]. Design [?]. Detection [?, ?]. Deuterium [?]. Developed [?, ?, ?, ?]. Developing [?]. Development [?, ?, ?, ?, ?, ?, ?, ?]. Device [?]. Devices [?, ?, ?, ?, ?, ?]. Diabetes [?]. Diagnosis [?, ?]. Diamond [?, ?]. Did [?, ?]. Differentiation [?]. Difficult [?]. Digest [?]. Dimensional [?, ?]. Dimensions [?, ?, ?, ?]. Dinosaur [?]. Diode [?]. Diodes [?]. Dioxide [?]. Direct [?]. Directed [?]. Dirty [?]. Disarmament [?]. Disclinations [?]. Discover [?]. Discovery [?, ?, ?]. Discrimination [?]. Disease [?, ?, ?, ?, ?]. Diseases [?]. Disk [?]. Disorders [?]. Disoriented [?]. Display [?, ?]. Displays [?]. Disposal [?, ?]. Dissection [?]. Distributed [?]. Diversity [?]. DNA [?, ?, ?, ?, ?, ?, ?, ?]. Do [?]. Doctor [?]. Doctor-Patient [?]. does [?]. Dolphins [?]. dominoes [?]. Double [?, ?]. Drained [?]. Drawings [?]. Dried [?]. Drift [?, ?, ?]. Drip [?]. Drowned [?]. Drug [?, ?]. Drug-Delivery [?]. Drums [?]. Dual [?, ?]. Dual-Resonance [?, ?]. Duet [?]. Duet-Singing [?]. Dung [?, ?]. Dust [?]. Dwarfism [?]. Dying [?]. Dynamic [?]. Dynamics [?, ?].

Earlier [?]. Earliest [?, ?, ?]. Early [?, ?, ?, ?, ?]. Earth [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Earth-Moon [?]. Earthquake [?]. Earthquakes [?]. East [?]. Eastern [?]. Ecological [?]. Ecology [?]. Economic [?, ?, ?, ?]. Economy [?, ?, ?]. Ecosystem [?, ?, ?]. Eden [?]. Effect [?, ?]. Effects [?, ?]. Efficiency [?, ?]. Eggs [?, ?]. Eggshell [?]. Eggshells [?]. Elastic [?]. Elder [?]. Electric [?]. Electric-Power [?]. Electrical [?]. Electricity [?]. Electrochemical [?]. Electromagnetic [?]. Electron [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Electron-Hole [?, ?]. Electron-Positron [?, ?, ?, ?]. Electronic [?, ?, ?]. Electronics [?, ?, ?]. Electrons [?]. Electrostatics [?]. Elementary [?, ?, ?, ?]. Elementary-Particle [?]. Elements [?, ?, ?]. Embryo [?]. Embryos [?, ?]. Emission [?]. Emitting [?, ?]. Endemic [?]. Energetics [?, ?]. Energies [?, ?]. Energy [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Energy-Storage [?, ?]. Enforcing [?]. Engine [?, ?]. Engines [?]. England [?, ?]. Enhanced [?]. Enhanced-Radiation [?]. Environment [?, ?]. Environmental [?]. Enzyme [?]. Enzymes [?]. Epidemiology [?]. Epstein [?]. Eradication [?]. Escape [?]. Escher [?]. Ethics [?, ?]. Euphrates [?]. Eurasia [?]. Europe [?, ?]. European [?]. Evoked [?]. Evolution [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Excavation [?]. Exchange [?]. Exoelectrons [?]. Exotic [?, ?]. Expand [?]. Experience [?]. Experiment [?, ?]. Experimental [?]. Experimentation [?]. Experiments [?, ?, ?, ?, ?]. Exploring [?]. Expression [?]. Extending [?]. External [?]. Extraterrestrial [?]. Eye [?, ?, ?, ?]. Eye-Head [?]. Eyes [?]. Eyewitness [?].

Fabrication [?]. Faces [?]. Factor [?, ?]. Factors [?, ?]. Fall [?]. Families [?]. Family [?, ?]. Fantastic [?, ?]. Farmer [?]. Farming [?, ?]. Farmstead [?]. Fast [?, ?]. Faster [?]. Fastest [?]. Fat [?]. Fat-Metabolism [?]. Fault [?]. Feedback [?]. Feeding [?, ?]. Ferdinand [?]. Fermat [?]. Fertility [?, ?]. Fertilization [?, ?]. Fiber [?]. Fibers [?]. Fibroplasia [?]. Field [?]. Fields [?]. Figures [?]. Films [?]. Final [?, ?]. Fine [?]. Finite [?]. Fire [?]. Fireflies [?]. First [?, ?, ?]. Fishes [?, ?, ?, ?]. Fission [?, ?, ?, ?, ?]. Fission-Track [?, ?]. Fixation [?, ?]. Flashlight [?]. Flavor [?]. Flea [?]. Flew [?]. Flight [?, ?, ?, ?]. Flight-Control [?]. Flint [?, ?]. Floor [?, ?, ?]. Flow [?, ?, ?, ?, ?, ?]. fluctuations [?]. Fluorescence [?, ?]. Fluorescence-Activated [?, ?]. Fly [?]. Flying [?, ?, ?]. Flywheels [?]. Food [?, ?, ?, ?, ?, ?]. Food-sharing [?]. Footracing [?]. Forage [?]. Force [?]. Forces [?]. Foreign [?, ?]. Forest [?, ?, ?]. Forever [?]. Form [?]. Formation [?, ?, ?, ?]. Forming [?]. Fortification [?]. Fortress [?]. Fossil [?, ?]. Fossils [?]. Four [?]. Four-Color-Map [?]. fractal [?]. Free [?, ?, ?]. Freedom [?]. Frequency [?]. Frog [?]. Frontier [?]. Fruit [?]. Fuel [?, ?, ?]. Fuel-Cell [?, ?]. Fuels [?, ?]. Full [?, ?]. Full-Scale [?, ?]. Function [?, ?, ?]. Functional [?]. Functions [?]. Fundamental [?, ?]. Fusion [?, ?, ?, ?, ?, ?, ?]. Future [?].

Galaxies [?, ?, ?, ?, ?, ?, ?, ?, ?]. Galaxy [?, ?, ?]. Galilean [?]. Galileo [?, ?, ?, ?]. Game [?, ?, ?, ?]. Games [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Games [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Gamma [?, ?]. Gamma-Ray [?, ?]. Garden [?]. Gas [?, ?, ?, ?]. Gaseous [?]. Gasification [?]. Gauss [?]. Gears [?]. Gene [?, ?]. Generation [?, ?]. Genes [?, ?, ?, ?, ?]. Genetic [?, ?, ?, ?, ?, ?]. Genetics [?, ?, ?]. Geography [?]. Geometry [?]. Geosynclines [?]. Geothermal [?]. Getting [?, ?, ?]. Giant [?, ?, ?]. Gigantopithecus [?]. Giordano [?]. Giotto [?]. Giraffe [?]. Giving [?]. Gland [?]. Global [?, ?]. Globular [?, ?, ?, ?]. Globular-Cluster [?, ?]. Globules [?]. Glory [?]. Glycoproteins [?]. go [?]. Goiter [?]. Gothic [?]. Government [?]. Graaff [?]. Gradient [?, ?]. Graduate [?]. Grape [?]. Graphics [?]. Gravitation [?]. Gravitational [?]. Gravity [?, ?]. Grazing [?]. Great [?, ?, ?]. Greece [?]. Greek [?, ?]. Green [?]. Ground [?, ?]. Grouse [?]. Growing [?, ?]. Growth [?, ?, ?, ?]. Guests [?]. Guide [?]. Guided [?, ?, ?]. Guided-Wave [?, ?]. Gum [?].

Hadrons [?]. Hailstones [?]. Hales [?]. Halley [?]. Hallucinations [?]. Halos [?]. Happens [?]. Head [?, ?]. Health [?, ?]. Heartbeat [?]. Heat [?]. Heavy [?, ?, ?]. Heavy-Water [?, ?]. Helium [?, ?, ?]. Hemoglobin [?, ?]. Hepatitis [?]. Herbarium [?]. Hereditary [?]. Herpes [?]. High [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. High-Efficiency [?]. High-Energy [?, ?, ?]. High-Gradient [?, ?]. High-Lysine [?]. High-Resolution [?, ?]. Highest [?, ?]. Highest-Energy [?, ?]. Hilbert [?]. Hilltop [?]. Histocompatibility [?]. Historical [?]. History [?, ?, ?, ?, ?]. Hole [?, ?]. Holes [?, ?, ?]. Holograms [?, ?]. Homing [?]. Hominid [?]. Hominids [?, ?]. Hormonal [?]. Hormones [?, ?, ?]. Hospital [?]. Host [?]. Hot [?, ?]. Housework [?]. Human [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Hunger [?]. Hunters [?]. Hunting [?]. Hybrid [?]. Hydra [?]. Hydraulic [?]. Hydrogen [?, ?, ?]. Hyperactive [?]. Hypodermic [?].

Hypothalamus [?]. hypothesis [?].

Ice [?]. Ice-Age [?]. Ideology [?]. II [?]. Ills [?]. Illusions [?, ?, ?]. Image [?, ?]. Images [?]. Imaginary [?]. Immune [?, ?, ?]. Immunization [?]. Immunology [?, ?]. Implantable [?]. Implantation [?]. Implosion [?]. Importation [?]. Imprinting [?]. Improvement [?]. Inapparent [?]. Income [?]. Index [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. India [?, ?]. Indian [?, ?, ?, ?]. Indicators [?]. Individuality [?]. Indochina [?]. Induced [?]. Induction [?, ?]. Industrial [?]. Inertial [?]. Infant [?, ?, ?, ?]. Infants [?]. Infectious [?]. Inflation [?]. Influence [?]. Influenza [?]. Information [?, ?]. Infrared [?, ?]. Ink [?, ?]. Ink-Jet [?, ?]. Inorganic [?]. Insect [?]. Insects [?, ?, ?, ?]. Insert [?]. instrumentation [?]. Integrated [?]. Integration [?, ?, ?, ?, ?, ?]. Intelligence [?, ?, ?]. Interaction [?]. Interactions [?, ?, ?]. Interactive [?]. Intercellular [?]. Intercontinental [?, ?]. Interference [?]. Interferon [?, ?]. Intergroup [?]. Interior [?, ?]. Internal [?]. International [?, ?, ?]. Interplanetary [?]. Interstellar [?, ?]. Intervention [?, ?, ?]. Intravenous [?]. Intrinsically [?]. Invertebrates [?]. Investment [?]. Ion [?, ?]. Iran [?]. Iranian [?]. Iron [?, ?, ?]. Iroquois [?]. Irrigation [?, ?]. Isolation [?]. Isotopes [?]. Issue [?]. itself [?].

Japan [?]. Japanese [?]. Jenner [?]. Jet [?, ?]. Job [?]. John [?]. Joint [?]. Journal [?]. Judaism [?]. jump [?]. Junctions [?]. Jupiter [?, ?, ?].

Kangaroos [?]. Karate [?]. Kepler [?]. Key [?]. Kilauea [?]. Kimberlite [?]. kind [?]. Kittens [?]. Kiwi [?]. Knee [?]. Knots [?].

Labor [?]. Laboratory [?, ?]. Labrador [?]. Lacertae [?]. Lactase [?]. Lactose [?]. Lake [?]. Lake-Bottom [?]. Lakes [?]. Land [?, ?]. Land-Based [?]. Language [?, ?, ?, ?, ?]. Languages [?, ?]. Large [?, ?, ?, ?]. Large-Scale [?, ?, ?, ?]. Laser [?, ?, ?, ?, ?, ?, ?, ?]. Lasers [?, ?, ?]. Last [?]. Lava [?]. Law [?]. Laws [?, ?]. Layers [?]. Lead [?]. Leap [?]. Learning [?, ?, ?]. Lectins [?]. Leech [?]. Left [?]. Legal [?]. Legionellosis [?]. Lek [?]. Length [?]. Lens [?, ?]. Leonardo [?]. Leptons [?]. Lesson [?]. Letters [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Letters [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Lewis [?]. Life [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Lift [?]. Light [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Light-Wave [?, ?]. Limitation [?, ?]. Limited [?]. Line [?]. Linear [?]. Linear-Chain [?]. Lines [?]. Lions [?]. Liquefied [?]. Liquid [?, ?, ?, ?]. Liquid-Crystal [?, ?]. Liquids [?, ?]. Lithosphere [?, ?, ?]. Liver [?]. Lives [?]. Living [?, ?]. Lizard [?]. Load [?]. Locusts [?]. Logic [?]. Long [?]. Long-Wavelength [?]. Looking [?, ?]. Lost [?]. Love [?]. loving [?]. Low [?, ?]. Lunar [?, ?, ?]. Lung [?]. Lymphocyte [?]. Lysine [?].

Macaques [?]. Machining [?, ?]. made [?]. Magnetic [?, ?, ?, ?, ?]. Magnetism [?]. Magnets [?]. Majorities [?]. Make [?]. making [?]. Malaria [?]. Malpractice [?]. Mammals [?, ?]. Man [?, ?, ?, ?, ?, ?, ?, ?]. Man-Carrying [?]. Man-Day [?]. Managed [?, ?]. Manipulation [?]. Mantle [?, ?]. Manufacture [?, ?]. many [?]. Map [?]. Mapping [?]. Marine [?, ?]. Mariner [?]. Markers [?]. Market [?]. Marriage [?, ?]. Mars [?, ?, ?, ?, ?, ?, ?]. Masers [?]. Mass [?, ?]. Massive [?]. Materials [?, ?, ?, ?, ?, ?]. Mathematical [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Mathematical [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Mathematics [?, ?, ?]. Mathemtical [?]. Mating [?]. Matrixes [?]. Matter [?, ?]. Maurits [?]. Maya [?, ?, ?]. Measurement [?]. Mechanical [?, ?, ?]. Mechanics [?]. Mechanism [?]. Mechanisms [?, ?, ?, ?, ?, ?]. Medical [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Medication [?]. Medicine [?, ?]. Medieval [?, ?]. Meditation [?]. Mediterranean [?]. Medium [?]. Melanism [?]. Membrane [?, ?, ?, ?]. Membranes [?, ?, ?, ?]. Memories [?]. Memory [?, ?, ?, ?, ?]. Mendel [?]. Mental [?]. Merchant [?]. Mercury [?, ?]. Mesozoa [?]. messenger [?]. Messengers [?]. Metabolism [?, ?]. Metabolizes [?]. Metal [?, ?, ?, ?, ?, ?]. Metal-Oxide [?]. Metal-Vapor [?, ?]. Metals [?, ?, ?, ?]. Metastasis [?]. Meteorites [?, ?]. Meteorology [?]. Metric [?]. Mexico [?]. Micro [?]. Micro-electronics [?]. Microbial [?]. Microcircuits [?, ?]. Microcomputers [?]. Microelectronic [?, ?, ?, ?, ?]. Microelectronics [?, ?, ?, ?, ?, ?]. Microprocessors [?]. Microscope [?, ?, ?, ?]. Microstructure [?, ?]. Microvascular [?]. Microwave [?]. Mid [?, ?]. Mid-Atlantic [?, ?]. Middle [?]. Migraines [?]. Migration [?]. Migrations [?, ?]. Migratory [?]. Military [?]. Millikan [?]. Millimeter [?]. millions [?]. Mimicry [?]. Mind [?]. Mine [?]. Mineral [?, ?]. Mining [?]. Mirages [?]. Mirror [?]. Missile [?]. Missiles [?, ?, ?, ?]. Missing [?]. Mississippi [?]. Mitochondria [?]. Model [?, ?, ?]. Models [?, ?, ?, ?]. Molecular [?, ?, ?]. Molecule [?]. Molecules [?, ?, ?, ?]. Mollusks [?]. Monomolecular [?]. monoxide [?]. Moon [?, ?, ?, ?]. Mosquito [?]. Mosquitoes [?]. Mössbauer [?]. Most [?, ?]. Mother [?]. Moths [?, ?, ?]. Motion [?, ?, ?]. Motions [?]. Mound [?]. Mountains [?]. Move [?, ?]. Movement [?, ?, ?]. Movements [?, ?, ?]. Moving [?, ?, ?]. Mule [?]. Multicellular [?]. Multiple [?, ?]. Multiple-Warhead [?]. Multistability [?]. Muscle [?, ?, ?]. muscles [?]. Muscular [?]. Mushrooms [?]. Music [?, ?]. Musical [?, ?]. Mutations [?]. Mycenaean [?]. Mystery [?].

National [?, ?]. Natural [?, ?, ?, ?, ?]. Natural-Uranium [?, ?]. Nature [?, ?, ?, ?, ?]. Navigation [?, ?, ?]. Neanderthals [?]. Nebula [?, ?]. Nebulas [?, ?]. Necessity [?]. Negative [?, ?, ?, ?]. Negative-Income-Tax [?]. Negotiations [?]. Neoglaciation [?]. Neolithic [?, ?, ?]. Nerve [?, ?, ?, ?, ?, ?]. Nerve-Cell [?, ?]. Nerve-Growth [?]. Nervous [?, ?, ?]. Network [?]. Networks [?, ?]. Neural [?]. Neurobiology [?]. Neuron [?]. Neurons [?]. Neurophysiology [?]. Neurotransmitters [?]. Neutral [?, ?]. Neutrino [?]. Neutron [?, ?]. Neutron-scattering [?]. Neutrons [?]. Newborn [?, ?]. Newfoundland [?]. Nighttime [?]. Nile [?, ?]. Nitrogen [?, ?, ?]. Nonsocial [?]. Nonstandard [?]. Nonvisual [?]. Nourish [?]. Nuclear [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Nuclear-Free [?]. Nuclear-Test [?]. Nuclei [?]. Nucleotide [?]. Number [?, ?]. Numbers [?]. Nutrient [?]. Nutrition [?, ?, ?].

Object [?]. Objects [?, ?, ?, ?]. Ocean [?, ?, ?, ?]. Oceanic [?, ?, ?, ?, ?]. Oceans [?, ?]. Off [?]. Offensive [?]. Oil [?, ?]. Old [?]. Oldest [?, ?]. Omnivorous [?]. Opals [?]. Operator [?]. Operator-Repressor [?]. Opiate [?]. Opiates [?]. Optical [?, ?]. Optics [?, ?, ?, ?, ?]. Oracle [?]. Order [?, ?]. Ore [?]. Organic [?, ?]. Organisms [?]. Organization [?, ?, ?]. Orientation [?, ?]. Origin [?, ?, ?, ?, ?, ?, ?]. Origins [?, ?, ?, ?, ?, ?, ?, ?]. Other [?]. Otherwise [?]. our [?]. Outer [?]. Oxide [?]. Oxygen [?].

Pacific [?]. Paleolithic [?, ?]. Paleoneurology [?]. Pangaea [?]. Parabolic [?]. Paradox [?]. Parasite [?]. Parasitic [?]. parsec [?]. Particle [?, ?, ?, ?]. Particle-Beam [?]. Particles [?, ?, ?, ?, ?, ?, ?, ?, ?]. Parts [?, ?]. Passive [?]. Pathways [?, ?]. Patient [?]. Pattern [?, ?, ?]. patterns [?]. Patterson [?]. Peer [?]. Pelagic [?]. People [?]. Perception [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Performance [?, ?]. Permanent [?]. Pernor [?]. Personal [?]. Personality [?]. Peru [?]. Pesticides [?]. Phenomena [?]. Phobos [?]. Phosphenes [?]. Photographic [?]. Photon [?]. Photons [?]. Photosynthesis [?, ?]. Photosynthetic [?]. Photovoltaic [?]. Physical [?]. Physics [?, ?, ?, ?, ?]. Physiological [?]. Physiology [?, ?, ?, ?, ?]. Piano [?]. Pictorial [?]. Pieter [?]. Pigeon [?]. Pigments [?]. Pineal [?]. Pines [?]. Pipes [?]. Pivot [?]. Placebos [?]. Plain [?]. Plan [?]. Planets [?, ?]. Planning [?]. Plant [?, ?]. Plants [?, ?, ?, ?, ?, ?, ?]. Plate [?, ?, ?]. Plays [?]. Pluralistic [?]. Poetic [?]. Poisoning [?]. Poisonous [?]. Poker [?]. Polarized [?]. Polarized-Light [?]. Policy [?, ?]. Poliovirus [?]. Political [?]. Pollutants [?]. Polycyclic [?]. Polymeric [?]. Polymers [?]. Pons [?]. Population [?, ?, ?, ?, ?, ?, ?, ?]. Populations [?, ?, ?, ?]. Portrait [?]. Positron [?, ?, ?, ?]. Positrons [?]. Post [?]. Potential [?]. Power [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Powered [?]. Pre [?, ?]. Pre-Columbian [?]. Pre-Neolithic [?]. Precursor [?]. Prediction [?]. Prehistoric [?]. Prehistory [?, ?]. Prematurity [?]. Prenatal [?]. Preservation [?]. Pressure [?, ?, ?, ?, ?]. Primeval [?]. Primitive [?]. Principle [?]. Printing [?, ?]. Prints [?]. Privacy [?]. Probe [?]. Problem [?, ?, ?, ?, ?, ?]. Problems [?, ?, ?]. Process [?, ?, ?, ?]. Processes [?]. Processing [?]. Product [?]. Production [?, ?, ?, ?, ?, ?, ?]. Professionalization [?]. Program [?]. programmed [?]. Programming [?]. Progress [?]. Projections [?]. Proliferation [?]. Prompt [?]. Proof [?, ?]. Prospects [?, ?]. Prostaglandins [?]. Protection [?]. Protein [?, ?, ?, ?]. Protein-Cutting [?]. Proteins [?, ?, ?]. Protohuman [?]. Proton [?, ?, ?, ?]. Psychiatric [?]. Psychiatrists [?]. Psychological [?]. Public [?, ?]. Public-Key [?]. Pulp [?]. Pulsars [?]. Pulsating [?]. Pupfish [?]. Pupil [?]. Purple [?].

Quantum [?, ?]. Quark [?]. Quarks [?, ?]. Quasars [?]. Quasi [?]. Quasi-Stellar [?]. Question [?]. Qumran [?].

Rabbit [?]. Race [?, ?]. Racial [?, ?]. Radar [?, ?]. Radiation [?, ?, ?, ?, ?]. Radicals [?]. Radio [?, ?]. radioactive [?]. Rain [?, ?]. Rainbow [?]. Ramapithecus [?]. Randomness [?]. Rangelands [?]. Rat [?]. Ray [?, ?, ?, ?, ?, ?, ?, ?, ?]. Rays [?, ?, ?, ?, ?]. Razed [?]. Reactions [?, ?]. Reactor [?, ?, ?, ?]. Reactors [?, ?, ?, ?]. Reading [?]. Real [?]. Real-Number [?]. Reality [?]. Reception [?]. Receptor [?]. Receptors [?, ?, ?]. Recognition [?, ?, ?]. Recombinant [?]. Recombinant-DNA [?]. Reconnaissance [?]. Reconstruction [?]. Record [?]. Recreations [?]. Recurrent [?]. Red [?, ?]. Reduction [?]. Reefs [?, ?]. Reflector [?]. Reflectors [?]. Regeneration [?]. Regulation [?]. Relations [?]. Remember [?]. Remnants [?, ?, ?]. Renaissance [?]. Repeated [?]. Repel [?]. Repellents [?]. Repetitive [?]. Replacement [?]. Repressor [?]. Repressors [?]. Reprocessing [?]. Reproduction [?, ?, ?]. Requirements [?]. Research [?, ?]. Reservoirs [?]. Resistance [?, ?]. Resolution [?, ?]. Resonance [?, ?]. Resources [?, ?, ?, ?, ?]. Respiratory [?]. Response [?, ?, ?]. Responses [?, ?, ?]. Retina [?]. Retinex [?]. Retrolental [?]. Review [?]. Revolution [?, ?]. Reward [?]. Ribosome [?]. Rich [?]. Rickets [?]. Rift [?, ?]. Right [?]. Rings [?]. Rise [?, ?]. River [?]. RNA [?, ?]. RNA-Directed [?]. Robert [?]. Robot [?]. Rocks [?, ?]. Rodents [?]. Role [?, ?, ?, ?, ?, ?, ?]. Roman [?, ?, ?, ?]. Roots [?]. Rotating [?]. Rotation [?, ?, ?, ?]. row [?]. Rules [?]. Running [?]. Rural [?].

Saccadic [?]. Sacred [?]. Sage [?]. Salt [?, ?, ?, ?, ?]. Salt-loving [?]. Samaritans [?]. San [?]. Sand [?]. Sardinia [?]. Satellite [?]. Satellites [?]. Satisfaction [?]. Saving [?]. Scale [?, ?, ?, ?, ?, ?]. Scales [?]. Scanning [?, ?, ?]. scattering [?]. Scheduling [?]. School [?]. Science [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Science [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Science-Textbook [?]. Scientific [?, ?]. Scientist [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Sclerosis [?]. Sea [?, ?, ?, ?, ?, ?]. Search [?, ?, ?, ?]. Sebei [?]. Second [?, ?, ?, ?]. Secretion [?]. Security [?, ?]. See [?]. Segments [?]. Seine [?]. Self [?]. Self-Reproduction [?]. Semiconductor [?, ?, ?]. Sensing [?]. Sensitivity [?]. Separation [?, ?, ?]. Sequence [?]. Serengeti [?]. Services [?]. Settlements [?]. Sex [?]. Sex-Attractant [?]. Shad [?]. Shale [?]. Shape [?, ?, ?, ?]. Shape-Memory [?, ?]. Shapes [?, ?]. Shaping [?]. sharing [?]. Sheet [?]. Short [?]. Short-Term [?]. Sickle [?]. Sickle-Cell [?]. Side [?, ?]. Side-Looking [?, ?]. Silent [?]. Silviculture [?]. similarity [?]. Singing [?, ?]. Sister [?]. Sister-Exchange [?]. Site [?, ?]. Size [?]. Sky [?, ?]. Slavery [?]. Sleep [?]. Sling [?]. Slips [?]. Slow [?]. Small [?, ?]. Smaller [?]. Smallpox [?, ?]. Snakes [?, ?]. Snow [?, ?]. Soap [?]. Soaring [?]. Social [?, ?, ?, ?, ?, ?, ?, ?]. Societies [?]. Society [?, ?, ?]. Soil [?]. Solar [?, ?, ?, ?, ?, ?, ?, ?, ?]. Solid [?, ?, ?, ?]. Solidification [?, ?]. Solids [?]. Solitaire [?]. Solitons [?]. Solution [?]. some [?]. Song [?, ?]. Sorting [?, ?]. Sound [?, ?]. Source [?, ?]. Sources [?, ?, ?]. Southern [?, ?, ?]. Soybeans [?]. Space [?, ?]. Spatial [?, ?]. Special [?]. Specializations [?]. Specificity [?]. Spectroscopy [?, ?]. Spectrum [?, ?]. Speech [?, ?]. Spending [?]. Spent [?]. Sperm [?]. Spiders [?]. Spin [?]. Spots [?]. Spread [?]. Stability [?]. Stairs [?]. Standard [?]. Star [?]. Stars [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Starvation [?]. State [?, ?, ?]. Stationary [?]. Statistics [?]. Status [?, ?]. Steady [?]. Stein [?]. Stellar [?, ?, ?]. Stellar-Orientation [?]. Stephen [?]. Steps [?]. Steroid [?]. Stethoscope [?]. Stick [?]. Stirling [?]. Stomach [?]. Stomatopods [?]. Stone [?, ?]. Stone-Age [?]. Storage [?, ?]. Storms [?]. Strategic [?, ?]. Strategy [?, ?]. Strength [?]. Stress [?, ?]. String [?]. Strings [?]. Strip [?]. Strip-Mining [?]. Stroke [?]. Structural [?]. Structure [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Studies [?]. Subduction [?]. Subjective [?]. Subjects [?]. Submarines [?]. Substances [?]. Succession [?]. Successors [?]. Sulfur [?]. Summer [?]. Sun [?, ?]. Sunlike [?]. Sunspots [?]. Superconductivity [?, ?]. Superconductors [?, ?, ?]. Superdense [?]. Superfluid [?]. Supergravity [?]. Superhard [?, ?]. Superior [?]. Supernova [?, ?, ?, ?]. Supernovas [?, ?, ?]. Superphenix [?, ?]. Support [?]. Surface [?, ?, ?, ?, ?, ?, ?, ?, ?]. Surgery [?, ?]. Surgical [?, ?]. Swim [?]. Switch [?]. Symbiosis [?]. Synchronous [?]. Synchrotron [?]. Synthesis [?, ?, ?, ?]. Synthetic [?]. Synthetic-Membrane [?]. System [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Systems [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?].

take [?]. Taking [?, ?]. Talking [?]. Tall [?]. Tandem [?]. Tar [?]. Targets [?, ?]. Task [?]. Tax [?]. Teaching [?]. Technology [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Tectonics [?, ?, ?]. Tektite [?]. Telephone [?]. Television [?]. Telling [?]. Temperature [?]. Temperatures [?, ?, ?]. Temple [?, ?]. Tensile [?]. Tephra [?]. Teratomas [?]. Term [?]. Terminals [?]. Test [?, ?]. Test-Ban [?]. Testimony [?]. Tests [?]. Textbook [?]. Texture [?, ?]. Their [?, ?]. Theorem [?]. Theories [?]. Theory [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Therapeutic [?]. thermodynamics [?]. Thermostat [?]. Think [?]. Thinking [?]. Three [?, ?, ?]. Three-dimensional [?]. Thunder [?]. Tidal [?]. Tides [?, ?, ?]. Tillage [?]. Time [?, ?, ?, ?]. Timekeeping [?]. Tissue [?]. Tissues [?]. Tokamak [?, ?, ?]. Tomato [?]. Tongue [?]. Tools [?]. Top [?]. toppling [?]. Total [?]. Towers [?]. traced [?]. Track [?, ?]. Tracks [?, ?]. Trading [?]. Trajectories [?]. Trajectory [?]. Transfer [?, ?]. Transmission [?]. Transparency [?]. Transplant [?]. Transplantation [?]. Transport [?]. Tree [?]. Trees [?, ?, ?]. Tremor [?]. Trends [?]. Triangle [?]. trigger [?]. Triticale [?]. Tropical [?]. Tube [?]. Tumors [?]. Turkana [?]. Turkey [?]. Turkeys [?]. Turned [?]. Two [?, ?]. Two-Dimensional [?]. Tycho [?].

U.S. [?, ?, ?, ?, ?]. Ukraine [?]. Ultracold [?]. Ultrafast [?]. Ultrasound [?]. Underdeveloped [?, ?]. Underground [?]. Unification [?]. Unified [?]. Uniqueness [?]. Universe [?, ?, ?, ?]. Unsolved [?]. Unusual [?]. Upsilon [?]. Uranium [?, ?]. Urban [?, ?]. Uses [?, ?].

Vacuoles [?]. Vacuum [?]. Vapor [?, ?]. Vascularization [?]. Venture [?]. Venus [?, ?]. Verbal [?]. Verification [?]. Vertebrate [?]. Vietnam [?]. Views [?]. Village [?, ?, ?, ?]. Vines [?]. Violate [?]. Violent [?, ?]. Viral [?, ?]. Virus [?, ?, ?]. Viruses [?, ?, ?]. Viscosity [?]. Vision [?, ?, ?]. Visual [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Visualization [?]. Visually [?]. Voice [?]. Volcanoes [?]. Vortexes [?]. Voting [?]. Vultures [?].

Wakes [?]. Walking [?]. Walls [?]. Wankel [?]. War [?, ?, ?]. Warfare [?]. Warhead [?]. Warm [?]. was [?]. Wasps [?]. Waste [?]. wastes [?]. Water [?, ?, ?, ?, ?, ?, ?, ?]. Wave [?, ?, ?, ?]. Wavelength [?]. Waves [?, ?, ?]. Wax [?]. Ways [?]. Weak [?]. Weaker [?, ?]. Weapon [?]. Weapons [?, ?, ?, ?, ?, ?, ?]. Wear [?]. Weaver [?]. Wegener [?]. Western [?, ?]. Whale [?]. Whistles [?]. White [?, ?, ?]. White-Light [?, ?]. Whites [?]. Will [?]. Winchester [?]. Wind [?, ?]. Winds [?]. Wines [?]. Winning [?]. without [?, ?]. Women [?, ?]. Wood [?].

Woodhenges [?]. Words [?]. Work [?]. World [?, ?, ?, ?, ?, ?]. Worm [?, ?]. would [?]. Wright [?]. Writing [?].

X [?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. X-ray [?, ?, ?, ?, ?]. X-ray-emitting [?, ?]. X-rays [?].

Yarn [?]. Year [?]. Years [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. Years [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?]. York [?].

Zone [?]. Zones [?]. Zoo [?].